### v1

# FERMILAB-PUB-25-0130-PPD

Abstract
Understanding electron neutrino interactions is crucial for measurements of neutrino oscillations and searches for new physics in neutrino experiments. We present the first measurement of the flux-averaged and charged current single charged pion production cross section on argon using the MicroBooNE detector and data from the NuMI neutrino beam. The total cross section is measured to be (0.93 0.13 (stat.) 0.27 (syst.)) cm2/nucleon at a mean and energy of 730 MeV. Differential cross sections are also reported in electron energy, electron and pion angles, and electron-pion opening angle.
Introduction.— The next generation of accelerator neutrino oscillation experiments will seek to address multiple open questions in neutrino physics through precision measurement of electron neutrino appearance in muon neutrino beams. These include the presence and scale of charge-parity violation in the neutrino sector, the neutrino mass hierarchy, and the resolution of anomalies observed at short baselines [1]. Several of these experiments will make use of the liquid argon time projection chamber (LArTPC) detector technology [2, 3]. As such, in order to facilitate these measurements, precise understanding of interactions on argon is essential.
Cross section modeling of interactions is typically extrapolated from measurements. However, uncertainties on the interaction cross section ratio limit their constraining power [4, 5, 6, 7]. Understanding the nuclear effects unique to -argon interactions is crucial for precision neutrino oscillation measurements, rare-event searches, and benchmarking theoretical models used in neutrino experiments.
Reconstructing and measuring interactions also presents unique challenges compared to interactions due to the electromagnetic cascades produced by final-state electrons. These are complicated to reconstruct and sensitive to detector calibration uncertainties. Measurements of interactions allow algorithms targeting these topologies to be evaluated and improved reducing systematic uncertainties and enhancing detector performance.
Existing measurements of cross sections on argon are limited consisting of several inclusive measurements [8, 9, 10] and a measurement without final state pions [11]. The Deep Underground Neutrino Experiment (DUNE) will be exposed to a neutrino flux peaking at energies of a few GeV [3]. At these energies, one of the dominant neutrino interaction modes leads to the excitation of baryon resonances that subsequently decay producing pions. This process has never previously been measured for interactions.
This work presents the first measurement of the flux-averaged and charged-current (CC) single charged pion production cross section on argon using MicroBooNE. The final-state topology considered consists of an electron (or positron), a single charged pion, zero neutral pions (or heavier mesons), and any number of protons or neutrons:
where represents the residual nucleus and any outgoing nucleons. These interactions will subsequently be referred to as for simplicity.
The MicroBooNE detector is an 85 tonne LArTPC that collected data between 2015 and 2020. It consists of an instrumented argon volume of () m3 (drift, vertical, beam direction). Ionization charge produced by charged particles resulting from neutrino interactions is drifted towards three planes of readout wires, orientated vertically and at to the vertical, by an electric field of 273 V/cm. Additionally, scintillation light is collected by an array of 32 photomultiplier tubes located behind the readout planes [12].
MicroBooNE collected data from two neutrino beams: the on-axis 8 GeV Booster Neutrino Beam and the approximately off-axis 120 GeV Neutrinos at the Main Injector (NuMI) beam [13]. In this work, data from the NuMI beam operating in a mixture of forward-horn-current (FHC) neutrino mode and reverse-horn-current (RHC) anti-neutrino mode are used. The NuMI flux at MicroBooNE contains a significant fraction of (%) due to the high beam energy and the significantly off-axis position of the detector. This makes it particularly useful for cross-section measurements. The average energies incident on MicroBooNE are MeV in FHC mode and MeV in RHC mode. The integrated exposure is protons-on-target (POT) in FHC mode and POT in RHC mode. This corresponds to the full NuMI dataset accumulated by MicroBooNE and used for the first time in this analysis.
At these neutrino energies, the dominant production mechanism is through the resonance that subsequently decays to a pion and a nucleon. In addition, there are subdominant contributions from coherent pion production and pion production as a result of deep inelastic scattering. The observed topology is also impacted by final-state interactions that can lead to pion production or absorption as the particles produced in the neutrino interaction leave the argon nucleus.
Although pion production has not previously been measured in interactions on argon, it has been probed in interactions. ArgoNeuT has performed measurements of and induced charged pion production on argon [14]. MicroBooNE has performed several measurements of induced neutral pion production on argon for both CC [15, 16] and neutral-current (NC) interactions [17, 18]. This work presents the first measurement of on argon. The total interaction cross section is reported along with differential cross sections in electron energy, electron and pion angles, and electron-pion opening angle. These measurements are complementary to the existing suite of measurements.
Simulation and reconstruction.— The NuMI neutrino flux is simulated with GEANT4 v4.10.4 [19, 20, 21, 22] constrained with available hadron production data using the PPFX software package [23]. The integrated flux of with energy over 60 MeV is /cm2/POT in FHC mode and /cm2/POT in RHC mode. The combined integrated flux is calculated by weighting the contributions from each horn-current mode according to the accumulated POT in that mode.
The LArSoft software framework [24] is used to perform simulation and reconstruction. Neutrino interactions are modeled using the GENIE v3.0.6 G18_10a_02_11a event generator [25] with the MicroBooNE tune applied [26]. In particular, resonant pion production is simulated using the Kuzmin-Lyubushkin-Naumov Berger-Sehgal model [27, 28, 29, 30] and coherent pion production using the Berger-Sehgal model [31]. Propagation of the final state particles through the detector is then performed using GEANT4 [19, 20, 21]. This is followed by simulation of the produced ionization electrons and scintillation light and the subsequent detector response [32, 33]. Simulated neutrino interactions are overlaid onto data collected while the beam is off providing data-based modeling of cosmic-ray induced interactions and detector noise. The Monte-Carlo (MC) prediction consists of simulated neutrino interactions within the detector, simulated interactions upstream of the detector (out-of-cryostat), and data collected with the beam off to model beam spills where no neutrino interaction occurs (EXT).
Reconstruction is performed using the Pandora pattern-recognition toolkit [34]. This uses a multi-algorithm approach to identify neutrino interactions from cosmic-ray-induced backgrounds and to reconstruct each resulting particle. Each particle associated with the neutrino interaction is categorized as a track (muons, pions, protons) or an electromagnetic shower (electrons, photons). Particle identification based on calorimetric and topological information is then performed [35, 36]. Finally, energy reconstruction is performed using particle range for tracks and through calorimetry for showers.
Signal, selection and observables.— Signal events are defined as or CC interactions that contain an outgoing electron or positron with kinetic energy KE MeV, a single charged pion with KE MeV, zero neutral pions or heavier mesons, and any number of outgoing protons or neutrons. The signal definition thresholds are guided by the reconstruction thresholds for each particle type. In addition, the opening-angle between the electron and charged pion is required to be . This removes a region of phase-space, containing less than 0.5% of predicted signal events, for which the reconstruction performance is poor. Figure 1 shows a candidate interaction matching the signal definition that is selected in MicroBooNE data. A high energy electromagnetic shower is visible along with a single track that is consistent with a charged pion.
The event selection expands upon tools developed in previous MicroBooNE analyses [9, 10, 37, 11]. Charged current interactions are first identified through the presence of an electromagnetic shower consistent with an electron. Tracks present in the event are then assumed to have originated from either protons or charged pions. These are distinguished based on their differing ionization profiles or the presence of pion re-interactions. The focus on first identifying charged current interactions avoids the need to distinguish between final state muons and charged pions that are otherwise more challenging to separate.
Well-reconstructed candidate events are first identified that have at least one electromagnetic shower and at least one track contained within the instrumented volume. Next, a set of simple cuts are used to remove obvious background events. Cosmic-ray-induced interactions are rejected through Pandora classification based on topological characteristics. Electromagnetic showers resulting from neutral pion decays to photons are rejected by applying cuts on three features: the fraction of energy in the leading shower compared to all showers; the shower start position relative to the interaction vertex; and the transverse spread of the shower. Finally, interactions containing protons but not pions (referred to subsequently as CC N) are rejected by requiring at least one track has an ionization profile () that is inconsistent with a stopping proton Bragg peak hypothesis [38].
After these initial cuts, more sophisticated methods are applied to refine the event selection further focusing on background suppression through the use of two Boosted Decision Trees (BDTs) trained with XGBoost [39]. The BDTs are trained separately for FHC and RHC modes accounting for differences in neutrino composition and energies.
The first BDT focuses on the shower information to distinguish between electrons and photons from neutral pion backgrounds ( CC , CC and NC ). It uses both calorimetric and topological information about the primary shower, any secondary showers if present, and any shower-like clusters of charge identified close to the neutrino interaction.
The second BDT focuses on the track information and aims to distinguish between charged pions and protons, in particular targeting remaining CC N interactions where the proton does not have a clear Bragg peak and, hence, is not rejected at the previous stage. This BDT uses the of the track compared to different particle hypotheses along with topological information focusing on the ends of tracks to identify pion re-interactions or decays. Interactions are selected where only one track is identified as a charged pion candidate, with any number of additional proton candidates allowed.
Figure 2 shows the distribution of scores for the -rejection BDT and CC N-rejection BDT compared with data. The CC N-rejection BDT is shown after the cut on the -rejection BDT has been applied. Both are shown for FHC mode; RHC mode can be found in the Supplemental Material [40]. Both BDTs achieve good signal-background separation for their target topologies. In addition, excellent data–MC agreement within statistical and systematic uncertainties is seen across the full BDT score distributions.
Candidate CC interactions are selected with efficiencies of 6.2% and 5.8% and purities of 57.3% and 58.5% for the FHC and RHC beam periods, respectively. In total, 116 candidate events pass the selection in the combined POT data sample. The selected signal events are predicted to be 78% interactions and 22% interactions. They are dominated by resonant production (85%), with smaller contributions from deep-inelastic scattering (11%), quasi-elastic interactions (2%) and coherent interactions (2%).
The observables considered are: the total rate; the electron energy, ; the electron and pion angles, and , with respect to the neutrino direction; and the electron-pion opening angle, . Since MicroBooNE is off-axis to the NuMI beam, the exact neutrino direction is not known in data. To reconstruct angles relative to the neutrino direction, neutrinos are assumed to be produced at the NuMI target position. This gives an effective approximation of the neutrino direction with 95% of signal events having an approximated direction within of the true direction. In simulation, the true direction is known and hence the smearing arising from this approximation can be accounted for during unfolding [10]. Details on the performance of this approximation can be found in the Supplemental Material [40]. The reconstruction resolution is approximately 17% for , 10% for , 10% for and 9% for .
Each observable is binned such that there are five bins with approximately ten expected signal events each. In the case of , the highest energy bin also serves as an overflow bin. For each observable, good data–MC agreement is observed within uncertainties. The selected event distributions can be found in the Supplemental Material [40]
The largest backgrounds are interactions producing neutral pions (13.0% of passing events) and CC N interactions (11.7% of passing events). Two sidebands are constructed to assess the agreement between data and simulation for these backgrounds. A -rich sideband is constructed by reversing the -rejection BDT cut. This results in a sample with 68.8% purity of -containing interactions with a mixture of CC , CC and NC topologies. A CC N-rich sideband is constructed by reversing both the CC N-rejection BDT cut and the proton Bragg peak cut. This results in a sample with 77.6% purity of CC N interactions. The level of agreement with data is assessed across each observable considered for both sidebands. Good data–MC agreement is seen across all distributions indicating the background modeling is sufficient to proceed with cross section extraction. The sideband selected event distributions can be found in the Supplemental Material [40].
Cross section extraction and uncertainties.— The flux-averaged total cross section and differential cross sections as a function of true kinematic variables are reported. The cross sections are extracted using the Wiener singular value decomposition unfolding technique [41] using a first-derivative regularization term. The impact of the regularization is encoded in a regularization matrix that can be applied to generator predictions to allow direct comparison with the extracted cross sections in the regularized truth space. A block-wise approach to unfolding is used allowing correlations between bins in different variables to be evaluated and reported [42].
Uncertainties on the extracted cross sections are assessed from a variety of sources. The statistical and systematic uncertainties are encoded in a covariance matrix using a block-wise formalism [42]. The total covariance matrix is constructed by summing the covariance matrices of each individual uncertainty.
Systematic uncertainties are considered on: the neutrino flux from hadron production and beam-line geometry modeling [23, 22]; the neutrino interaction cross section modeling with GENIE [26]; secondary particle re-interactions [43]; detector response modeling including the scintillation light yield, recombination model, space charge effects [44], and ionization signal response [45]; out-of-cryostat interaction modeling; the number of argon targets; and POT counting. The dominant systematic uncertainty arises from the neutrino flux modeling (20–30%) due to the challenges in simulating the off-axis NuMI beam. This is followed by detector response modeling (15%), predominantly from recombination modeling due to the reliance on calorimetric variables to identify charged pions; and cross section modeling (10%). The other sources of systematic uncertainty are subdominant.
Statistical uncertainties on the data and simulation are evaluated using Poisson uncertainties. Data statistical uncertainties are subdominant for the total cross-section measurement at around 10% but are comparable with systematic uncertainties or dominant for the differential cross-section measurements at around 30%. The total covariance and correlation matrices are reported in the Supplemental Material [40].
The robustness of the unfolding procedure and regularization is assessed using fake data produced by the GENIE and NuWro 19.02.2 [46] generator models. These fake-data tests motivated expanding the cross section modeling uncertainty by treating the NuWro sample as an additional systematic universe [47].
Results.— The extracted total cross section is shown in Table 1 and differential cross sections in , , and are shown in Fig. 3. The differential cross sections are presented in regularized truth space described by the regularization matrix available in the Supplemental Material [40]. The unfolded data is compared with generator predictions from NuWro 21.09.2 [46], NEUT 5.4.0.1 [48], GiBUU 2025 [49], GENIE 3.0.6 G18_10a_02_11a MicroBooNE tune [26] (labeled GENIE 3.0.6 G18B), and GENIE 3.4.2 AR23_20i_00_000 [25] (labeled GENIE 3.4.2 AR23). Resonant pion production is simulated using the Kuzmin-Lyubushkin-Naumov Berger-Sehgal model [27, 28, 29, 30] in GENIE, the Adler-Rarita-Schwinger formalism [50] in NuWro, and the Rein-Sehgal model in NEUT [51]. Coherent pion production is simulated with the Berger-Sehgal [31] model in GENIE, NuWro and NEUT. GiBUU models resonant pion production following the MAID analysis [52] and does not simulate coherent pion production.
| Generator | [cm2/nucleon] | |
| Unfolded Data | 0.93 0.13 (stat.) 0.27 (syst.) | |
| NuWro 21.09.2 | 0.76 | 0.30/1 |
| NEUT 5.4.0.1 | 0.83 | 0.11/1 |
| GiBUU 2025 | 0.74 | 0.42/1 |
| GENIE 3.4.0 AR23 | 0.62 | 1.11/1 |
| GENIE 3.0.6 G18 B | 0.68 | 0.68/1 |
The total cross section is measured to be (0.93 0.13 (stat.) 0.27 (syst.)) cm2/nucleon. This is consistent with the predictions from each of the generators considered. The data slightly prefers the higher cross sections predicted by NEUT, NuWro and GiBUU compared with the lower cross sections predicted by GENIE. This could suggest a slight preference for the treatment of resonant pion production or nuclear medium effects in these models. However, the sensitivity of the measurement is limited by the large systematic uncertainties on the flux modeling.
The extracted differential cross sections in , , and are all consistent with the generator predictions. The preference for higher cross sections results in lower for NEUT, NuWro and GiBUU in , and ; whereas these models have a higher for hinting at possible shape disagreement in this variable. The largest are seen for driven by the smallest opening angle bin where, in particular, GENIE and GiBUU underpredict the cross section. However, due to the large statistical and flux uncertainties, all models lie within or close to 1 of the data suggesting in general good modeling of this process within the sensitivity of this measurement. To achieve greater model separation, future measurements would require greater statistics, potentially through improved reconstruction, and improved flux modeling. In particular, additional hadron production data to constrain the off-axis NuMI flux could significantly reduce the associated uncertainties.
Conclusions.— We have presented the first measurement of the flux-averaged charged current single charged pion production cross section on argon using the MicroBooNE detector. The full NuMI beam data set accumulated by MicroBooNE is used for the first time corresponding to a total exposure of POT with mean energy of 730 MeV. The total cross section is measured to be (0.93 0.13 (stat.) 0.27 (syst.)) cm2/nucleon. Differential cross sections are also reported as functions of electron energy, the electron and pion angles with respect to the neutrino direction, and the electron-pion opening angle. These are found to be in good agreement with generator predictions within uncertainties. This is the first measurement of pion production in interactions, one of the dominant interaction modes at the energies of the DUNE neutrino flux. It is essential to understand this process in order to facilitate the DUNE physics program, and direct measurements will have a critical impact to motivate modeling improvements. While this is the first time this process has been measured, the measurement is limited by statistical uncertainties and the challenges arising from the off-axis flux; future experiments in the SBN program [2] and the DUNE near detector [53] are expected to be able to improve on this.
Acknowledgments.— This document was prepared by the MicroBooNE collaboration using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by Fermi Forward Discovery Group, LLC, acting under Contract No. 89243024CSC000002. MicroBooNE is supported by the following: the U.S. Department of Energy, Office of Science, Offices of High Energy Physics and Nuclear Physics; the U.S. National Science Foundation; the Swiss National Science Foundation; the Science and Technology Facilities Council (STFC), part of the United Kingdom Research and Innovation; the Royal Society (United Kingdom); the UK Research and Innovation (UKRI) Future Leaders Fellowship; and the NSF AI Institute for Artificial Intelligence and Fundamental Interactions. Additional support for the laser calibration system and cosmic ray tagger was provided by the Albert Einstein Center for Fundamental Physics, Bern, Switzerland. We also acknowledge the contributions of technical and scientific staff to the design, construction, and operation of the MicroBooNE detector as well as the contributions of past collaborators to the development of MicroBooNE analyses, without whom this work would not have been possible. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) public copyright license to any Author Accepted Manuscript version arising from this submission.